OpenStudy (babynini):

If a ball is thrown into the air with a velocity of 40 ft/s, its height (in feet) after t seconds is given by y = 40t − 16t^2. Find the velocity when t = 1. help!

OpenStudy (babynini):

@zepdrix

jimthompson5910 (jim_thompson5910):

the velocity is going to be equal to dy/dt

jimthompson5910 (jim_thompson5910):

apply the derivative and then plug in t = 1

OpenStudy (babynini):

Don't know derivative yet :/

OpenStudy (babynini):

Unless..I knw to use lim (h approaches 0) [f(a+h)-f(a)]/h

jimthompson5910 (jim_thompson5910):

ok we can use that

jimthompson5910 (jim_thompson5910):

f(t) = 40t - 16t^2 f(t+h) = ???

OpenStudy (babynini):

f(t+h) = 40(1+h)-16(1+h)^2

jimthompson5910 (jim_thompson5910):

use t instead of 1

jimthompson5910 (jim_thompson5910):

well I guess you can plug in t = 1 later, so it works out

OpenStudy (babynini):

For now just keep it t?

jimthompson5910 (jim_thompson5910):

I would keep t for now if I were doing it, but again you can plug in t = 1 when doing the limit definition

jimthompson5910 (jim_thompson5910):

so f(a+h) - f(a) = [ 40(1+h)-16(1+h)^2 ] - [ 24 ] = ??

OpenStudy (babynini):

40+40h-16+32h+16h^2-24

jimthompson5910 (jim_thompson5910):

-16(1+h)^2 = -16(1+2h+h^2) -16(1+h)^2 = -16-32h-16h^2

OpenStudy (babynini):

ooh dangit, that negative in front of the 16. Ah k.

OpenStudy (babynini):

Which simplifies to: 8h-16h^2

OpenStudy (babynini):

And further to: 8-16h once we factor out an h and cancel it.

OpenStudy (babynini):

then plugging in 0 we get 8 :)

jimthompson5910 (jim_thompson5910):

very nice, 8 is the correct answer

OpenStudy (babynini):

Thank you!

jimthompson5910 (jim_thompson5910):

you're welcome